A projectile is fired horizontally from a height of 10 m above level ground. The projectile lands a horizontal distance of 15 m from where it was launched.
1
Expert's answer
2020-12-10T11:02:14-0500

Answer

Vertical displacement can be given as

"10=\frac{1}{2}gt^2\\t=\sqrt{\frac{20}{9.8}}=1.4s"

Horizontal displacement can be given as

"15=v_0t\\v_0=\frac{15}{1.4}=10.7m/s"

So initial horizontal velocity is 10.7m/s